If the RBA surprises the market by raising its cash rate target to 3% the AUD will immediately __________ (appreciate / depreciate) against all other currencies. This immediate change of the AUD value is making our exports __________ (cheaper / more expensive) and thus __________ (fosters / dampens) economic activity and inflation. In Japan the interest rate is close to 0%. Therefore, the AUD is very likely to ____________ (appreciate / depreciate) against the Yen over the coming year. In Indonesia the interest rate is around 6%. Therefore, the AUD is very likely to _____________ (appreciate / depreciate) against the Indonesian Ruphia over the coming year.
Answer #1

If RBA increases the cash rate by 3% then the AUD will immediately appreciate against all other currencies.

Note: Generally a higher interest rate increases the value of domestic currency. A higher interest attract foreign investors which increases the value of domestic currency.

The immediate change of the AUD value making our exports more expensive thus dampens economic activity and Inflation.

In Japan the interest rate is 0%. Therefore, the AUD is very likely to depreciate against the Yen over the coming year.

In Indonesia the interest rate is around 6%. Therefore, the AUD is very likely to appreciate  against the Indonesian ruphia over the coming year.

Use the following formula

\large i_{AUD} -i_{Yen} = \frac{F - S}{S}

When s = AUD 1 per yen

Then, 0.03 - 0 = (F -1)/1

F = AUD 1.03 per yen

So depreciated against yen.

Similarly we can solve for Indonesia.

Please contact if having any query thank you .

IAUD-1 yen-

0.03 - 0.00 -